The lifetime of a machine is continuous on the interval (0, 40) with probability density function f, where f(t) is proportional to (t + 10)−2, and t is the lifetime in years. Calculate the probability that the lifetime of the machine part is less than 10 years. Hint: Show that f(t) is legitimate and find the proportionality constant.
1
Expert's answer
2020-05-12T18:59:51-0400

"f(t)=\alpha \frac{1}{(t+10)^2}", "t \in [0,40]" for some real "\alpha."


"f(t)" is density, so "\int_{-\infty}^{+\infty} f(t)dt =1".


Find "\alpha":

"\int_{-\infty}^{+\infty} f(t) dt= \alpha\int_{0}^{40} (t+10)^{-2}dt=\{ -\alpha (t+10)^{-1}\} \bigg|^{40}_{0}"

"=(-\alpha(40+10)^{-1}) - (-\alpha(0+10)^{-1})=-\alpha/50 + \alpha/10=2\alpha/25"

"2\alpha/25=1 => \alpha=25/2."

Probability the lifetime between 0 and 10 years is "\int_{0}^{10} f(t)dt.\\""\\ \text{Then} \int_{0}^{10} f(t)dt = \alpha\int_{0}^{10} (t+10)^{-2} dt"

"=25/2\cdot\int_0^{10}(t+10)^{-2}dt=25/2\cdot \{-(t+10)^{-1}\}\bigg|_{0}^{10}=25/2\cdot(-\frac{1}{20}-(-\frac{1}{10}))=\frac{5}{8}."


Answer: "\frac{5}{8}."